# The long way of a viscous vortex dipole

### v1

@@ 1. Introduction
where $z_{1}(t),\dots,z_{N}(t)\in\mathbb{R}^{2}$ denote the positions of the vortex centers and $\Gamma_{1},\dots,\Gamma_{N}\in\mathbb{R}$ are the circulations of the corresponding vortices. The ODE system (1.1), which was already studied by Helmholtz and Kirchhoff, can be rigorously derived as an asymptotic model for the evolution of sharply concentrated solutions of the two-dimensional Euler or Navier-Stokes equations [22, 21, 12, 6]. Note that (1.1) only makes sense if $z_{i}(t)\neq z_{j}(t)$ when $i\neq j$, but this condition is not always preserved under evolution. Indeed, if $N\geq 3$ and if the circulations $\Gamma_{i}$ do not have the same sign, there are examples of solutions of (1.1) for which three vortices collide in finite time [22]. However, the set of initial data leading to such collisions is negligible in the sense of Lebesgue’s measure.
@@
$\frac{1}{|\Gamma|}\int_{\mathbb{R}^{2}}\Bigl{|}\,\omega^{\nu}(x,t)\,-\,\sum_{i=1}^{N}\frac{\Gamma_{i}}{4\pi\nu t}\,e^{-\frac{|x-z_{i}(t)|^{2}}{4\nu t}}\Bigr{|}\,{\rm d}x\,\leq\,C_{0}\,\frac{\nu t}{d^{2}}\,,\qquad\text{for all}~{}\,t\in(0,T)\,,$ (1.7)
@@
Interesting questions arise when trying to extend the approximation result in Theorem 1.1 to longer time scales. The first one is related to the collapse of point vortices. Assume for instance that the solution of system (1.1) is defined on the maximal time interval $[0,T_{*})$, with three vortices collapsing at the origin as $t\to T_{*}$. In that case estimate (1.7) is valid on the time interval $(0,T)$ for any $T<T_{*}$, but the constant $C_{0}>0$ blows up as $T\to T_{*}$ because the distance $d$ converges to zero in this limit. So it is not clear at all if the weak convergence (1.8) holds up to collision time, although this is certainly a reasonable conjecture. A fortiori, we do not know if a limiting procedure as in (1.8) can be used to define a continuation of the point vortex dynamics after collapse.
In a different direction, one may consider global solutions of the point vortex system, for which $|z_{i}(t)-z_{j}(t)|\geq d>0$ for all $t\geq 0$ if $i\neq j$. This is the case, for instance, if $N=2$ or if the circulations $\Gamma_{i}$ are all positive [22]. Here again the approximation result (1.7) can be applied on any time interval $(0,T)$, but the constant $C_{0}$ depends on $T$ and increases rapidly as $T\to+\infty$. In particular, it is not clear if the solution $\omega^{\nu}(x,t)$ is well approximated by a superposition of Lamb-Oseen vortices as long as $\nu t\ll d^{2}$, namely when the radius of the vortex cores is small compared to the distance $d$ between the centers.
@@
so that $\omega^{in}$ represents a pair of point vortices of circulations ±$\pm\Gamma$ separated by a distance $d$. In that situation, the point vortex dynamics predicts a uniform translation at speed $V=\Gamma/(2\pi d)$ in a direction normal to the line segment joining the vortex centers. More precisely, the positions $z_{\ell}(t),z_{\mathrm{r}}(t)$ of the point vortices with circulation $\Gamma$, $-\Gamma$ (respectively) are given by
@@
$\omega_{\mathrm{app}}^{\nu}(x,t)\,=\,\frac{\Gamma}{4\pi\nu t}\Bigl{(}e^{-\frac{|x-z_{\ell}(t)|^{2}}{4\nu t}}-e^{-\frac{|x-z_{\mathrm{r}}(t)|^{2}}{4\nu t}}\Bigr{)}\,,\qquad x\in\mathbb{R}^{2}\,,\quad t>0\,,$ (1.11)
provided $\nu T\ll C_{0}^{-1}d^{2}$ where the constant $C_{0}$ depends on $T$.
@@
Our main result shows that the solution of (1.2) can be approximated by a viscous vortex dipole over a long time interval of size $T_{\mathsf{adv}}\delta^{-\sigma}$, where $10<\sigma<$ 1, provided that the position $Z_{2}(t)$ of the vortex centers (1.10) is chosen in a suitable way.
@@
Fix $\sigma\in[0,1)$. There exist positive constants $C$ and $\delta_{0}$ such that, for any $\Gamma>0$, any $d>0$ and any $\nu>0$ such that $\delta:=\nu/\Gamma\leq\delta_{0}$, the solution $\omega^{\nu}$ of (1.2) with initial data (1.9) satisfies
$\frac{1}{\Gamma}\int_{\mathbb{R}^{2}}\bigl{|}\,\omega^{\nu}(x,t)-\omega_{\mathrm{app}}^{\nu}(x,t)\bigr{|}\,{\rm d}x\,\leq\,C\,\frac{\nu t}{d^{2}}\,,\qquad\text{for all}~{}\,t\in\bigl{(}0,T_{\mathsf{adv}}\delta^{-\sigma}\bigr{)}\,,$ (1.13)
where $\omega_{\mathrm{app}}^{\nu}$ is the viscous vortex dipole (1.11), (1.10) and the vertical position $Z_{2}(t)$ of the vortex centers is a smooth function satisfying $Z_{2}(0)=0$ and
@@
The second step in the proof consists in showing that the exact solution of the vorticity equation (1.2) remains close to the approximation constructed in the first step. This is far from obvious because the linearized equation at the approximate solution contains very large advection terms that could potentially trigger violent instabilities, and also because we need a control over the long time interval $(0,T_{\mathsf{adv}}\delta^{-\sigma})$. We follow here the earlier study [14] which considers the related, but more complicated, problem of a an axisymmetric vortex ring in the low viscosity limit. The idea is to control the difference between the solution and its approximation using suitable energy functionals, which incorporate the necessary information about the stability of the approximate solution. Unlike in [12], it does not seem possible to rely on weighted enstrophy estimates only, so we follow the general approach introduced by Arnold to study the stability of steady states for the 2D Euler equations, see [13]. This method gives a clear roadmap to design a nonlocal energy-like functional whose evolution is barely affected by the most dangerous terms in the linearized equation at the approximate solution.
@@ 2.1. Self-similar variables
Here we write {,}$\{\cdot,\cdot\}_{x}$ or {,}$\{\cdot,\cdot\}_{\xi}$ according to whether the Poisson bracket is computed with respect to the original variable $x$ or the rescaled variable $\xi$. Recalling that $\delta=\nu/\Gamma$, we deduce from the equalities above that (1.2) is satisfied provided
@@
This is exactly the equation satisfied by the profile of a self-similar solution of the Navier-Stokes equation (1.2). It thus follows from [16, Prop. 1.3] that any solution of (2.14) that is integrable over $\mathbb{R}^{2}$ is necessarily a multiple of the Lamb-Oseen vortex
$\Omega_{0}(\xi)\,=\,G(\xi),\qquad\text{where}\quad G(\xi)\,=\,\frac{1}{4\pi}\,e^{-|\xi|^{2}/4}\,.$ (2.15)
@@
$U_{0}(\xi)\,=\,U^{G}(\xi)\,:=\,\frac{1}{2\pi}\frac{\xi^{\perp}}{|\xi|^{2}}\,\Bigl{(}1-e^{-|\xi|^{2}/4}\Bigr{)}\,,\qquad\Psi_{0}(\xi)\,=\,\frac{1}{4\pi}\mathrm{Ein}\bigl{(}|\xi|^{2}/4\bigr{)}-\frac{\gamma_{E}}{4\pi}\,,$ (2.16)
@@ 2.2. Formula for the vertical speed
In fact the first condition in (2.17) is fulfilled if $Z_{1}(t)=d/2$, which we already assumed in (2.3). The second condition requires a specific choice of the vertical position $Z_{2}(t)$, which we now describe.
@@
Let $\Omega$ be the solution of (2.9) with initial data (2.15), and $U=\nabla^{\perp}\Psi$ be the associated velocity field. If the vertical velocity is chosen so that
$Z_{2}^{\prime}(t)\,=\,-\frac{\Gamma}{\varepsilon d}\int_{\mathbb{R}^{2}}U_{2}\bigl{(}\widetilde{\xi}-\varepsilon^{-1}e_{1},t\bigr{)}\Omega(\xi,t)\,{\rm d}\xi\,,\quad t>0\,,$ (2.18)
@@
[…]
Here and in what follows we use the notation (1.4) for the moments of $\Omega$.
@@
[…]
@@ 3. The approximate solution
The choice of the integer $M$, which determines the accuracy of the approximate solution, depends on the intended purpose. The leading order deformation of the stream lines and of the level sets of vorticity is already obtained for $M=2$, whereas the first correction to the vertical speed $Z_{2}^{\prime}$ only appears when $M=5$, see Section 3.4. In particular, this means that $\zeta_{k}=0$ for $k=1,2,3$. As we shall see in Section 4, we need to take $M>3$ if we want to control the solution of (1.2) over a time interval $[0,T_{\mathsf{adv}}]$ that is independent of the viscosity parameter. More generally, if $10\leq\sigma<$ 1, we need $M>(3+\sigma)/(1-\sigma)$ to reach the time $T_{\mathsf{adv}}\delta^{-\sigma}$.
@@ 3.1. Function spaces and operators
which is a Hilbert space equipped with the scalar product
@@
1) If $g_{\varepsilon}\in\mathcal{S}_{*}(\mathbb{R}^{2})$ depends on a small parameter $\varepsilon>0$, we say that $g_{\varepsilon}\,=\,\mathcal{O}_{\mathcal{S}_{*}}\bigl{(}\varepsilon^{M}\bigr{)}$ if
$\forall\,\alpha\in\mathbb{N}^{2}~{}\exists\,C>0~{}\exists\,N\in\mathbb{N}\quad\text{such that}\quad|\partial^{\alpha}g_{\varepsilon}(\xi)|\leq C(1+|\xi|)^{N}\varepsilon^{M}~{}\,\forall\,\xi\in\mathbb{R}^{2}\,.$
@@
see for instance [15, Appendix A]. The one-dimensional kernel of $\mathcal{L}$ is spanned by the Gaussian function $G$ defined in (2.15), whereas the first-order derivatives $\partial_{1}G,\partial_{2}G$ span the two-dimensional eigenspace corresponding to the eigenvalue $-1/2$. More generally, the eigenvalue $-n/2$ has multiplicity $n+1$ and the eigenfunctions are Hermite functions of degree $n$.
@@
Another important operator, denoted by $\Lambda:D(\Lambda)\to\mathcal{Y}$, arises when linearizing the quadratic term in (1.2) at the Lamb-Oseen vortex. The operator is defined by
@@
[…]
@@
[…]
@@
which is justified in Section A.1. This leads to the formula
@@
Since $\Omega\in\mathcal{Z}$, we can replace the integral in $B_{R_{\varepsilon}}$ with the integral in $\mathbb{R}^{2}$ up to an error of size $\mathcal{O}_{\mathcal{S}_{*}}(\varepsilon^{N+1})$. Thus (3.18) is proved in view of (3.17). ∎
@@ 3.2. The second order approximation
where the operators $\mathcal{T}_{\varepsilon}$ and $\mathbf{T}_{\varepsilon}$ are defined in (2.10) and (2.13), respectively. The following statement is a particular case of the results established in [12, Section 3.1]. We give a short proof for the reader’s convenience.
@@
$\mathcal{R}_{0}(\xi)\,=\,\frac{1}{2\pi}\biggl{(}\frac{\eta^{\perp}}{|\eta|^{2}}-\frac{(\xi+\eta)^{\perp}}{|\xi+\eta|^{2}}+\frac{(\xi+\eta)^{\perp}}{|\xi+\eta|^{2}}\,e^{-|\xi+\eta|^{2}/4}\biggr{)}\cdot\nabla G(\xi)\,.$ (3.23)
@@
It is important to notice that, according to (3.21), the remainder $\mathcal{R}_{0}$ is already of order $\varepsilon^{2}$. As is clear from the above proof, the cancellation of the first order in $\varepsilon$ is due to the choice $\zeta_{\mathrm{app}}=1$, which in turn is equivalent to $Z_{2}^{\prime}=\Gamma/(2\pi d)$. In other words, in our approach the translation speed given the Helmholtz-Kirchhoff system can be recovered by minimizing the error of the naive approximation. This is a general feature of interacting vortices in the plane, see [12]. In the case of axisymmetric vortex rings, which share many similarities with counter-rotating vortex pairs, the remainder of the naive approximation is $\mathcal{O}(\varepsilon)$ and not better, even if the translation speed is chosen appropriately, see [14].
Since $\mathcal{R}_{0}=\mathcal{O}_{\mathcal{Z}}(\varepsilon^{2})$, the first nontrivial step in our construction is the second order approximation, corresponding to $M=2$, for which $\Omega_{\mathrm{app}}=\Omega_{0}+\varepsilon^{2}\Omega_{2}$ and $\Psi_{\mathrm{app}}=\Psi_{0}+\varepsilon^{2}\Psi_{2}$. Although this is not immediately obvious, we anticipate that $\zeta_{1}=0$, which means that there is no correction to the translation speed at this level of the approximation. To determine the unknown profile $\Omega_{2}$, the strategy is again to minimize the remainder (3.3), which takes the form
$\begin{split}\mathcal{R}_{2}\,&=\,\delta\varepsilon^{2}(\mathcal{L}-1)\Omega_{2}+\Bigl{\{}\Psi_{0}-\mathcal{T}_{\varepsilon}\Psi_{0}+\frac{\varepsilon\xi_{1}}{2\pi}+\varepsilon^{2}\Psi_{2}-\varepsilon^{2}\mathcal{T}_{\varepsilon}\Psi_{2}\,,\,\Omega_{0}+\varepsilon^{2}\Omega_{2}\Bigr{\}}\\ \,&=\,\delta\varepsilon^{2}(\mathcal{L}-1)\Omega_{2}+\mathcal{R}_{0}+\varepsilon^{2}\Lambda\Omega_{2}+\varepsilon^{2}\mathcal{N}_{2}\,,\end{split}$ (3.25)
@@
The properties recalled in Section 3.1 imply that the linear operator $\delta(\mathcal{L}-1)+\Lambda$ is maximal dissipative in the Hilbert space $\mathcal{Y}$, hence invertible for any $\delta>0$. Since $\mathcal{H}_{2}\in\mathcal{Z}\subset\mathcal{Y}$, there exists a unique profile $\Omega_{2}\in\mathcal{Y}$ such that the quantity inside brackets in (3.28) vanishes exactly. However, we need to verify that $\Omega_{2}\in\mathcal{Z}$ and that $\Omega_{2}$ does not blow up in the limit $\delta\to 0$, which is not immediately obvious. For these reasons, we find it simpler to solve the problem approximately, using only the information given by Proposition 3.6. We set $\Omega_{2}=\Omega_{2}^{E}+\delta\Omega_{2}^{NS}$, where
@@
It is possible to obtain a more explicit expression of the profile $\Omega_{2}$, in terms of solutions of linear ODEs on $\mathbb{R}_{+}$, see Section 3.4. In particular $\Omega_{2}^{E}(\xi)=(\xi_{2}^{2}-\xi_{1}^{2})w_{2}(|\xi|)$ for some smooth and nonnegative function $w_{2}$ with Gaussian decay at infinity. The profile $\Omega_{2}^{E}$, which represents the leading order correction to the radially symmetric vortex $\Omega_{0}$ in the expansion (3.1), is responsible for the deformation of the stream lines and of the level lines of the vorticity, which are nearly elliptical at this order of approximation, see Fig. 1.

### v2

@@ 1. Introduction
where $z_{1}(t),\dots,z_{N}(t)\in\mathbb{R}^{2}$ denote the positions of the vortex centers and $\Gamma_{1},\dots,\Gamma_{N}\in\mathbb{R}$ are the circulations of the corresponding vortices. The ODE system (1.1), which was already studied by Helmholtz and Kirchhoff, can be rigorously derived as an asymptotic model for the evolution of sharply concentrated solutions of the two-dimensional Euler or Navier-Stokes equations [24, 23, 13, 6]. Note that (1.1) only makes sense if $z_{i}(t)\neq z_{j}(t)$ when $i\neq j$, but this condition is not always preserved under evolution. Indeed, if $N\geq 3$ and if the circulations $\Gamma_{i}$ do not have the same sign, there are examples of solutions of (1.1) for which three vortices collide in finite time [18]. However, the set of initial data leading to such collisions is negligible in the sense of Lebesgue’s measure [24].
@@
$\frac{1}{|\Gamma|}\int_{\mathbb{R}^{2}}\Bigl|\,\omega^{\nu}(x,t)\,-\,\sum_{i=1}^{N}\frac{\Gamma_{i}}{4\pi\nu t}\,{\rm e}^{-\frac{|x-z_{i}(t)|^{2}}{4\nu t}}\Bigr|\,{\rm d}x\,\leq\,C_{0}\,\frac{\nu t}{d^{2}}\,,\qquad\text{for all}~\,t\in(0,T)\,$, (1.7)
@@
Interesting questions arise when trying to extend the approximation result in Theorem 1.1 to longer time scales. The first one is related to the collapse of point vortices. Assume for instance that the solution of system (1.1) is defined on the maximal time interval $0,T_{*})$, with three vortices collapsing at the origin as $t\to T_{*}$. In that case estimate (1.7) is valid on the time interval $(0,T)$ for any $T<T_{*}$, but the constant $C_{0}>0$ blows up as $T\to T_{*}$ because the distance $d$ converges to zero in this limit. So it is not clear at all if the weak convergence (1.8) holds up to collision time, although this is certainly a reasonable conjecture. A fortiori, we do not know if a limiting procedure as in (1.8) can be used to define a continuation of the point vortex dynamics after collapse.
In a different direction, one may consider global solutions of the point vortex system, for which $|z_{i}(t)-z_{j}(t)|\geq d>0$ for all $t\geq 0$ if $i\neq j$. This is the case, for instance, if $N=2$ or if the circulations $\Gamma_{i}$ are all positive [24]. Here again the approximation result (1.7) can be applied on any time interval $(0,T)$, but the constant $C_{0}$ depends on $T$ and increases rapidly as $T\to+\infty$. In particular, it is not clear if the solution $\omega^{\nu}(x,t)$ is well approximated by a superposition of Lamb-Oseen vortices as long as $\nu t\ll d^{2}$, namely whenever the radius of the vortex cores is small compared to the distance $d$ between the centers.
@@
so that $\omega^{in}$ represents a pair of point vortices of circulations $\pm\Gamma$ separated by a distance $d$. In that situation, the point vortex dynamics predicts a uniform translation at speed $V=\Gamma/(2\pi d)$ in a direction normal to the line segment joining the vortex centers. More precisely, the positions $z_{\ell}(t),z_{\mathrm{r}}(t)$ of the point vortices with circulation $\Gamma$, $\Gamma$ (respectively) are given by
@@
$\omega_{\mathrm{app}}^{\nu}(x,t)\,=\,\frac{\Gamma}{4\pi\nu t}\Bigl({\rm e}^{-\frac{|x-z_{\ell}(t)|^{2}}{4\nu t}}-{\rm e}^{-\frac{|x-z_{\mathrm{r}}(t)|^{2}}{4\nu t}}\Bigr)\,,\qquad x\in\mathbb{R}^{2}\,,\quad t>0\,$, (1.11)
provided $\nu T\ll C_{0}^{-1}d^{2}$ where the constant $C_{0}$ depends on $\Gamma T/d^{2}$.
@@
Our main result shows that the solution of (1.2) can be approximated by a viscous vortex dipole over a long time interval of size $T_{\mathsf{adv}}\delta^{-\sigma}$, where $0<\sigma<1$, provided that the position $Z_{2}(t)$ of the vortex centers (1.10) is chosen in a suitable way.
@@
Fix $\sigma\in[0,1)$. There exist positive constants $C$ and $\delta_{0}$ such that, for any $\Gamma>0$, any $d>0$ and any $\nu>0$ with $\delta:=\nu/\Gamma\leq\delta_{0}$, the solution $\omega^{\nu}$ of (1.2) with initial data (1.9) satisfies
$\frac{1}{\Gamma}\int_{\mathbb{R}^{2}}\bigl|\,\omega^{\nu}(x,t)-\omega_{\mathrm{app}}^{\nu}(x,t)\bigr|\,{\rm d}x\,\leq\,C\,\frac{\nu t}{d^{2}}\,,\qquad\text{for all}~\,t\in\bigl(0,T_{\mathsf{adv}}\delta^{-\sigma}\bigr)\,.$ (1.13)
Here $\omega_{\mathrm{app}}^{\nu}$ is the viscous vortex dipole (1.11), (1.10) and the vertical position $Z_{2}(t)$ of the vortex centers is a smooth function satisfying $Z_{2}(0)=0$ and
@@
The second step in the proof consists in showing that the exact solution of the vorticity equation (1.2) remains close to the approximation constructed in the first step. This is far from obvious because the linearized equation at the approximate solution contains very large advection terms that could potentially trigger violent instabilities, and also because we need a control over the long time interval $(0,T_{\mathsf{adv}}\delta^{-\sigma})$. We follow here the earlier study [15] which considers the related, but more complicated, problem of an axisymmetric vortex ring in the low viscosity limit. The idea is to control the difference between the solution and its approximation using suitable energy functionals, which incorporate the necessary information about the stability of the approximate solution. Unlike in [13], it does not seem possible to rely on weighted enstrophy estimates only, so we follow the general approach introduced by Arnold to study the stability of steady states for the 2D Euler equations, see [14]. This method gives a clear roadmap to design a nonlocal energy-like functional whose evolution is barely affected by the most dangerous terms in the linearized equation at the approximate solution.
@@
To conclude this introduction, we mention that the techniques developed in the present paper have recently been extended to study vortex pairs with general circulations [27], as well as more complicated configurations called vortex crystals [10].
@@ 2.1. Self-similar variables
Here we write $\{\cdot,\cdot\}_{x}$ or $\{\cdot,\cdot\}_{\xi}$ according to whether the Poisson bracket is computed with respect to the original variable $x$ or the rescaled variable $\xi$. Recalling that $\delta=\nu/\Gamma$, we deduce from the equalities above that (1.2) is satisfied provided
@@
This is exactly the equation satisfied by the profile of a self-similar solution of the Navier-Stokes equation (1.2). It thus follows from [17, Prop. 1.3] that any solution of (2.14) that is integrable over $\mathbb{R}^{2}$ is necessarily a multiple of the Lamb-Oseen vortex
$\Omega_{0}(\xi)\,=\,G(\xi),\qquad\text{where}\quad G(\xi)\,=\,\frac{1}{4\pi}\,{\rm e}^{-|\xi|^{2}/4}\,.$ (2.15)
@@
$U_{0}(\xi)\,=\,U^{G}(\xi)\,:=\,\frac{1}{2\pi}\frac{\xi^{\perp}}{|\xi|^{2}}\,\Bigl(1-{\rm e}^{-|\xi|^{2}/4}\Bigr)\,,\qquad\Psi_{0}(\xi)\,=\,\frac{1}{4\pi}\mathrm{Ein}\bigl(|\xi|^{2}/4\bigr)-\frac{\gamma_{E}}{4\pi}\,$, (2.16)
@@ 2.2. Formula for the vertical speed
In fact the first condition in (2.17) is fulfilled if $Z_{1}(t)=d/2$, which we already assumed in (2.3). The second condition requires a specific choice of the vertical position $Z_{2}(t)$, which we now describe. Here and in what follows we use the notation (1.4) for the moments of $\Omega$.
@@
Let $\Omega$ be the solution of (2.9) with initial data (2.15), and $U=\nabla^{\perp}\Psi$ be the associated velocity field. If for all $t>0$ the vertical velocity is chosen so that
$Z_{2}^{\prime}(t)\,=\,-\frac{\Gamma}{\varepsilon d}\int_{\mathbb{R}^{2}}U_{2}\bigl(\widetilde{\xi}-\varepsilon^{-1}e_{1},t\bigr)\Omega(\xi,t)\,{\rm d}\xi\,\equiv\,\frac{\Gamma}{\varepsilon d}\int_{\mathbb{R}^{2}}\bigl(\partial_{1}\mathcal{T}_{\varepsilon}\Psi\bigr)(\xi,t)\Omega(\xi,t)\,{\rm d}\xi\,$, (2.18)
@@
$\mathrm{M}[\Omega(t)]\,=\,1\,,\qquad\mathrm{m}_{1}[\Omega(t)]\,=\,\mathrm{m}_{2}[\Omega(t)]\,=\,0\,,\qquad t>0\,.$ (2.19)
@@
$=\,\frac{1}{2\pi}\iint_{\mathbb{R}^{4}}\frac{\xi_{2}-\eta_{2}}{|\xi_{1}+\varepsilon^{-1}+\eta_{1}|^{2}+|\xi_{2}-\eta_{2}|^{2}}\,\Omega(\xi)\,\Omega(\eta)\,{\rm d}\xi\,{\rm d}\eta\,=\,0\,$,
@@ 3. The approximate solution
The choice of the integer $M$, which determines the accuracy of the approximate solution, depends on the intended purpose. The leading order deformation of the stream lines and of the level sets of vorticity is already obtained for $M=2$, whereas the first correction to the vertical speed $Z_{2}^{\prime}$ only appears when $M=5$, see Section 3.4. In particular, this means that $\zeta_{k}=0$ for $k=1,2,3$. As we shall see in Section 4, we need to take $M>3$ if we want to control the solution of (1.2) over a time interval $0,T_{\mathsf{adv}}]$ that is independent of the viscosity parameter. More generally, if $0\leq\sigma<1$, we need $M>(3+\sigma)/(1-\sigma)$ to reach the time $T_{\mathsf{adv}}\delta^{-\sigma}$.
@@ 3.1. Function spaces and operators
which is a Hilbert space equipped with the inner product
@@
1) If $g_{\varepsilon}\in\mathcal{S}_{*}(\mathbb{R}^{2})$ depends on a small parameter $\varepsilon>0$, we say that $g_{\varepsilon}\,=\,\mathcal{O}_{\mathcal{S}_{*}}\bigl(\varepsilon^{M}\bigr)$ if, for all $\alpha\in\mathbb{N}^{2}$ and all $m\in\{0,\dots,M\}$, there exist $C>0$ such that
$|\partial_{\xi}^{\alpha}\partial_{\varepsilon}^{m}g_{\varepsilon}(\xi)|\,\leq\,C(1+|\xi|)^{N}\varepsilon^{M-m}\quad\forall\,\xi\in\mathbb{R}^{2}\,.$
@@
see for instance [16, Appendix A]. The one-dimensional kernel of $\mathcal{L}$ is spanned by the Gaussian function $G$ defined in (2.15), whereas the first-order derivatives $\partial_{1}G,\partial_{2}G$ span the two-dimensional eigenspace corresponding to the eigenvalue 1/2-1/2. More generally, the eigenvalue n/2-n/2 has multiplicity $n+1$ and the eigenfunctions are Hermite functions of degree $n$.
@@
Another important operator, denoted by $\Lambda:D(\Lambda)\to\mathcal{Y}$, arises when linearizing the quadratic term in (1.2) at the Lamb-Oseen vortex. It is defined by
@@
where $\Omega_{0},\Psi_{0}$ are also defined in (2.15), (2.16) and $\{\cdot,\cdot\}$ is the Poisson bracket (2.8). The operator $\Lambda$ is considered as acting on the maximal domain
@@
$\mathcal{P}_{0}\,g\,=\,0\,,\quad\text{and}\quad\mathrm{m}_{1}[g]\,=\,\mathrm{m}_{2}[g]\,=\,0\,$, (3.16)
@@
which is justified in Section A.1. Truncating the Taylor series at order $N$ and estimating the integral remainder together with its derivatives with respect to $\xi$ and $\varepsilon$, we obtain the formula
@@
Since $\Omega\in\mathcal{Z}$, we can replace the integral over $B_{R_{\varepsilon}}$ with the integral over $\mathbb{R}^{2}$ up to an error of size $\mathcal{O}_{\mathcal{S}_{*}}(\varepsilon^{N+1})$. Thus (3.18) is proved in view of (3.17). ∎
Remark 3.9.
In the sequel, when controlling the size of error terms, we shall use the notation $\mathcal{O}_{\mathcal{S}_{*}}(\varepsilon^{M})$ without explicitly checking that all conditions required by Definition 3.4 are satisfied. Indeed, derivatives with respect to $\xi$ are straightforward to estimate, and the dependence of our quantities with respect to $\varepsilon$ is either polynomial or occurs through the operator $\mathcal{T}_{\varepsilon}$. In the latter case, one can argue as in the proof of Lemma 3.8.
@@ 3.2. The second order approximation
where the operators $\mathcal{T}_{\varepsilon}$ and $\mathbf{T}_{\varepsilon}$ are defined in (2.10) and (2.13), respectively. The following statement is a particular case of the results established in [13, Section 3.1]. We give a short proof for the reader’s convenience.
@@
$\mathcal{R}_{0}(\xi)\,=\,\frac{1}{2\pi}\biggl(\frac{\eta^{\perp}}{|\eta|^{2}}-\frac{(\xi+\eta)^{\perp}}{|\xi+\eta|^{2}}+\frac{(\xi+\eta)^{\perp}}{|\xi+\eta|^{2}}\,{\rm e}^{-|\xi+\eta|^{2}/4}\biggr)\cdot\nabla G(\xi)\,.$ (3.23)
@@
It is important to notice that, according to (3.21), the remainder $\mathcal{R}_{0}$ is already of order $\varepsilon^{2}$. As is clear from the above proof, the cancellation of the first order in $\varepsilon$ is due to the choice $\zeta_{\mathrm{app}}=1$, which in turn is equivalent to $Z_{2}^{\prime}=\Gamma/(2\pi d)$. In other words, in our approach the translation speed given by the Helmholtz-Kirchhoff system can be recovered by minimizing the error of the naive approximation. This is a general feature of interacting vortices in the plane, see [13]. In the case of axisymmetric vortex rings, which share many similarities with counter-rotating vortex pairs, the remainder of the naive approximation is $\mathcal{O}(\varepsilon)$ and not better, even if the translation speed is chosen appropriately, see [15].
Since $\mathcal{R}_{0}=\mathcal{O}_{\mathcal{Z}}(\varepsilon^{2})$, the first nontrivial step in our construction is the second order approximation, corresponding to $M=2$, for which $\Omega_{\mathrm{app}}=\Omega_{0}+\varepsilon^{2}\Omega_{2}$ and $\Psi_{\mathrm{app}}=\Psi_{0}+\varepsilon^{2}\Psi_{2}$. Although this is not obvious, we anticipate that $\zeta_{1}=0$, which means that there is no correction to the translation speed at this level of approximation. To determine the unknown profile $\Omega_{2}$, the strategy is again to minimize the remainder (3.3), which takes the form
[…]
@@
The properties recalled in Section 3.1 imply that the linear operator $\delta(\mathcal{L}-1)+\Lambda$ is maximally dissipative in the Hilbert space $\mathcal{Y}$, hence invertible for any $\delta>0$. Since $\mathcal{H}_{2}\in\mathcal{Z}\subset\mathcal{Y}$, there exists a unique profile $\Omega_{2}\in\mathcal{Y}$ such that the quantity inside brackets in (3.28) vanishes exactly. However, we need to verify that $\Omega_{2}\in\mathcal{Z}$ and that $\Omega_{2}$ does not blow up in the limit $\delta\to 0$, which is not immediately obvious. For these reasons, we find it simpler to solve the problem approximately, using only the information given by Proposition 3.6. We set $\Omega_{2}=\Omega_{2}^{E}+\delta\Omega_{2}^{NS}$, where
@@
It is possible to obtain a more explicit expression of the profile $\Omega_{2}$, in terms of solutions of linear ODEs on $\mathbb{R}_{+}$, see Section 3.4. In particular $\Omega_{2}^{E}(\xi)=(\xi_{2}^{2}-\xi_{1}^{2})w_{2}(|\xi|)$ for some smooth and nonnegative function $w_{2}$ with Gaussian decay at infinity. The profile $\Omega_{2}^{E}$, which represents the leading order correction to the radially symmetric vortex $\Omega_{0}$ in the expansion (3.1), is responsible for the deformation of the stream lines and of the level lines of the vorticity, which are nearly elliptical at this order of approximation, see Fig. 1. As an aside, since $\mathcal{H}_{2}=\{h_{2},G\}$ for a suitable function $h_{2}$, it is observed in [10] that $\Omega_{2}^{E}$ can be represented by a Neumann series involving the operator $A^{-1/2}(-\Delta)^{-1}A^{-1/2}$, where $A$ is defined in (3.51) below. This is, in fact, a consequence of the analysis carried out in [14].